# Large Language Models Synergize with Automated Machine Learning

### v1

@@ A.1 Large Language Models
Iterative methods repeatedly generate the output sequences in order to improve the output sequence based on a sequence evaluation procedure. Let the initial input sequence be $X_{0}$, the sequence generation process be $f_{\lambda}$, and the evaluation procedure be $\mathbb{E}$. Then the output sequence in the $i_{th}$ iteration is $Y_{i}=f_{\lambda}(\Phi_{i}(Y_{<i},\mathbb{E}(Y_{<i}),X_{0}))$, where $\Phi_{i}$ represents a mechanism that assembles the input sequence for the iteration by integrating the output sequences, their evaluations, and the initial input sequence from preceding iterations. Recent works in this line include Yao et al. (2022); Madaan et al. (2023); Cobbe et al. (2021); Shinn et al. (2023); Chen et al. (2023b). In a reinforcement learning context, each output sequence represents an action Yao et al. (2022); Shinn et al. (2023). The implementation of our method incorporates the self-reflection mechanism in Shinn et al. (2023). More specifically, $\Phi_{i}$ in our implementation assembles input sequence from code execution results and an explanation of the results.
Prompting methods improves the output sequence $Y$ by transforming the input sequence $X$. Let $\psi$ represents the the prompt engineering process, then $p(Y|X)=\prod_{n=1}^{N}p_{\lambda}(y_{n}|y_{<n},\psi(X))$. Notable prompting methods including in-context learning Wei et al. (2022); Zhou et al. (2022a); Wang et al. (2022); Min et al. (2022); Yao et al. (2023) and zero-shot prompting Kojima et al. (2022).
@@ A.2 Automated Machine learning
Table 4 compares our method with existing methods that employ LLMs for autoML. A detailed analysis of recent advancements in this domain can be found in Tornede et al. (2023). To the best of our knowledge, Text-to-ML is the first method that automatically generates the optimized program of the entire ML pipeline across task modalities.
While recent autoML techniques also include predictor-based methods Dudziak et al. (2020) and weight-sharing methods Pham et al. (2018), these techniques are less suitable for the desiderata in our study: applications in low resource constraints scenarios and search space with diverse pretrained deep learning models.
Table 4: Comparison with existing methods in terms of qualitative features. (1) Program synthesis: whether the method automatically generates programs (2) Deep learning: whether the method automatically generates (part of) programs for deep learning tasks. (3) Data preparation: whether the method automatically chooses the strategy for data preparation (4) Modeling: whether the method automatically chooses the strategy for modeling (5) Optimization: whether the method optimizes the chosen components/pipelines. * Although the method can automatically choose the strategy for modeling, the method does not automatically generate the program for modeling. ** The method uses the LLM as a pseudo-optimization model instead of a dedicated optimization algorithm (e.g., evolutionary computation and reinforcement learning.)
@@ Methods
GPT-NAS Yu et al. (2023) $\checkmark$* $\checkmark$
MLCopilot Zhang et al. (2023a) $\checkmark$*
GENIUS Zheng et al. (2023) $\checkmark$* $\checkmark$**
Evoprompting Chen et al. (2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$**
CAAFE Hollmann et al. (2023) $\checkmark$ $\checkmark$
SapientML Saha et al. (2022) $\checkmark$ $\checkmark$ $\checkmark$
AutoML-GPT Zhang et al. (2023b) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ *
HuggingGPT Shen et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$
@@ Appendix B Proofs of Theorems
$=P(Y\in\Omega)$ (1)
Theorem. (Theorem 3.5) For any sequence generation process, if there exists a step in the process that generates the entire output sequence and Assumption 3.4 holds for the step, then the best case complexity of $\mathbf{E}(G)$ of the process is $\mathcal{O}(exp(\Gamma))$.
@@
In the best case, $\mathbf{p}_{X}^{\Omega,\lambda}(i)=\epsilon_{max}\;(1\leq i\leq I)$. Then by equation (1), $P(Y\in\Omega)=\epsilon_{max}^{I}$ and $\mathbf{E}(G)=\epsilon_{max}^{-I}$. Therefore, the best case complexity of $\mathbf{E}(G)$ is $\mathcal{O}(exp(\Gamma))$.
@@
Then, in the best case, $P(Y\in\Omega)=\epsilon^{\xi_{X}^{\Omega,\lambda}(\epsilon)}$ and $\mathbf{E}(G)=\epsilon^{-\xi_{X}^{\Omega,\lambda}(\epsilon)}$.
Theorem. (Theorem 3.8) If Assumption 3.7 holds, then for Contextual Modular Generation as defined in Algorithm 1, the best case complexity of $\mathbf{E}(G)$ is $\mathcal{O}(\Gamma)$.
@@
Then by equation (1), in the best case, for each $Y^{j}$, $P(Y^{j}\in\Omega^{j})=\epsilon_{min}^{Z^{j}}\geq\epsilon_{min}^{M}\;(1\leq j\leq J)$. In Contextual Modular Generation, each module is iteratively generated until being evaluated as valid. Therefore, we have
$\mathbf{E}(G)\leq\sum_{j=1}^{\lceil\frac{I}{M}\rceil}\mathbf{E}(G^{j})={\lceil\frac{I}{M}\rceil\epsilon_{min}^{-M}}$
Thus, the best case complexity of $\mathbf{E}(G)$ is $\mathcal{O}(\Gamma)$.
@@ C.1 Unit Testing Details
1:Input: a progenitor testing protocol $\mathfrak{P}$, an autoregressive LLM $f_{\lambda}$, the textual description of an ML task $\mathbb{T}$, a specification of the targeted module $\mathbb{M}$, and the number of versions of synthetic data $N$
@@
4:     Given $\mathbb{T}$ and $\mathfrak{P}$, $f_{\lambda}$ devises a plan $\mathbb{P}$ that details the specific features and data flow of the modules tailored for the task.
5:     Verify $\mathbb{P}$ with $\mathfrak{P}$
6:until $\mathbb{P}$ is allowed by $\mathfrak{P}$
7:With $\mathbb{P}$ and $\mathfrak{P}$, construct the specific unit tests for the ML task $\{\mathbb{E}^{i}\}$
@@
10:         Given $\mathbb{P}$, $f_{\lambda}$ generates program $\mathbf{D}_{n}$, which produces synthetic input and output data for the ML task
11:         Verify $\mathbf{D}_{n}$ with the test for synthetic data in $\{\mathbb{E}^{i}\}$
12:     until $\mathbf{D}_{n}$ passes the test in $\{\mathbb{E}^{i}\}$
@@
14:Given $\mathbb{T}$ and $\mathbb{M}$, $f_{\lambda}$ generates the program $\mathbf{C}_{\mathbb{M}}$ for $\mathbb{M}$
@@
16:     Verify $\mathbf{C}_{\mathbb{M}}$ with $\mathbf{D}_{n}$ and the test for the module in $\{\mathbb{E}^{i}\}$
17:     if $\mathbf{C}_{\mathbb{M}}$ passes the test then
@@
Regardless of the specific task and modules, the generated modules should be compatible with a pre-written set of modules for evaluation, training, and validation. Progenitor testing protocols first ensure the external connectivity by testing the form of the input and output data in the ML task, since all modules, including data preparation, modeling, post processing, optimizer, and loss functions, center around the input and output data in the ML task.
@@ C.3 Search Strategy
In our study, we employ two search strategies: random search and BOHB Falkner et al. (2018). Random search is a powerful baseline in hyperparameter optimization due to its wide applicability and search efficiency for tasks with low effective dimensionality Bergstra & Bengio (2012). BOHB combines Bayesian Optimization (BO) and Hyperband (HB) Li et al. (2018) to optimize hyperparameters efficiently. BOHB inherits HB’s robustness and flexibility, and BO’s ability to ensure strong performance. BOHB starts by rapidly identifying promising configurations like HB and refines them using BO to achieve optimal solutions faster. It’s designed to be simple, computationally efficient, and adaptable to various tasks. For BOHB, we use the default hyperparameters in Falkner et al. (2018) and set 30 epochs as the maximum budget for deep learning tasks.
@@ C.5 ZC proxies details
ZC proxies are an autoML technique that allows for the prediction of a fully trained neural network’s performance with small computational expenses. In this study, we utilize four ZC proxies: flops, params, naswot, and synflow. The term flops stands for Floating Point Operations Per Second, indicating the computational power of the network. Params denotes the total number of parameters within the network, reflecting its complexity. Although flops and params are fundamental concepts within deep learning, they have been validated as powerful baseline ZC proxies Krishnakumar et al. (2022).
@@ C.7 Hyperparameters for finetuning
Our research focuses on an unusual search space with pre-trained models at the architectural level. Öztürk et al. (2022) also explored a similar search space. Consequently, we utilize a comparable set of hyperparameters for fine-tuning the models (Table 5). To tailor this approach to our specific requirements, we have reduced the minimum batch size from 16 to 2 to accommodate the varying GPU memory demands across different tasks.
@@ D.2 Kaggle datasets
Age The ICR - Identifying Age-Related Conditions competition Kaggle & InVitro Cell Research (2023) involves the development of machine learning models capable of detecting various age-related conditions from medical imaging data. Participants are provided with a dataset containing images labeled with different age-related conditions, and the task is to create models that can accurately classify these conditions. The competition aims to advance research in medical imaging by leveraging AI to improve the diagnosis and understanding of age-related diseases.
@@
Default The American Express - Default Prediction competition Kaggle & Express (2023) challenges participants to predict the likelihood of credit default by American Express cardholders. The dataset provided includes a wide range of anonymized features spanning transactional data, account information, and user behavior over time, designed to simulate real-world conditions. The goal is to enhance credit risk models and improve financial inclusion by accurately identifying at-risk customers.
@@
Whale The Happywhale - Whale and Dolphin Identification competition Kaggle & Happywhale (2023) focuses on the identification of individual whales and dolphins from images. Participants are tasked with developing algorithms that can recognize and match cetaceans based on unique physical features captured in photographs, contributing to marine biology research and conservation efforts. The dataset comprises images labeled with species and individual IDs, making it a challenge in both classification and individual identification.
@@
Strip The Mayo Clinic - STRIP AI competition Kaggle & Clinic (2023) is aimed at predicting stroke and other thrombotic events by analyzing electroencephalogram (EEG) data. This competition provides a platform for the development of AI models that can assist clinicians in early detection and intervention for stroke patients, leveraging a dataset of EEG recordings with associated clinical outcomes. This challenge stands at the intersection of AI and healthcare, with the potential to significantly impact patient care and outcomes in neurology.
@@
Exam The Kaggle - LLM Science Exam competition Kaggle (2023) challenges participants to create language models that can solve science exam questions. The focus is on evaluating the ability of these models to understand and apply scientific knowledge accurately. This competition is part of Kaggle’s efforts to benchmark the performance of language models in educational contexts, providing a unique opportunity to advance AI in the field of automated question answering.
@@
Learning The Feedback Prize - English Language Learning competition Kaggle & Lab (2023) aims to develop AI models that can provide feedback on written English tasks by students learning English as a foreign language. The challenge involves analyzing student-written essays to identify areas for improvement, thus aiding in language learning and teaching. This competition seeks to leverage AI to enhance English language education by providing scalable, personalized feedback.
@@ Appendix E Prompts
Figure 14: Prompt for CIFAR-10. The loading code is copied from the official description of the datasets Krizhevsky et al. (2023)
Figure 15: Prompt for CIFAR-100. The loading code is copied from the official description of the datasets Krizhevsky et al. (2023)

### v2

@@ A.1 Large Language Models
Iterative methods repeatedly generate the output sequences in order to improve the output sequence based on a sequence evaluation procedure. Let the initial input sequence be $X_{0}$, the sequence generation process be $f_{\lambda}$, and the evaluation procedure be $\mathbb{E}$. Then the output sequence in the $i_{th}$ iteration is $Y_{i}=f_{\lambda}(\Phi_{i}(Y_{<i},\mathbb{E}(Y_{<i}),X_{0}))$, where $\Phi_{i}$ represents a mechanism that assembles the input sequence for the iteration by integrating the output sequences, their evaluations, and the initial input sequence from preceding iterations. Recent works in this line include Yao et al. (2022); Madaan et al. (2023); Cobbe et al. (2021); Shinn et al. (2023); Chen et al. (2023b). In a reinforcement learning context, each output sequence represents an action (Yao et al., 2022; Shinn et al., 2023). The implementation of our method incorporates the self-reflection mechanism in Shinn et al. (2023). More specifically, $\Phi_{i}$ in our implementation assembles input sequence from code execution results and an explanation of the results.
Prompting methods improves the output sequence $Y$ by transforming the input sequence $X$. Let $\psi$ represents the prompt engineering process, then $p(Y|X)=\prod_{n=1}^{N}p_{\lambda}(y_{n}|y_{<n},\psi(X))$. Notable prompting methods including in-context learning (Wei et al., 2022; Zhou et al., 2022a; Wang et al., 2022; Min et al., 2022; Yao et al., 2023; Li et al., 2023) and zero-shot prompting (Kojima et al., 2022). Notably, although some prompting methods break the problem into intermediate steps, such as Wei et al. (2022) and Li et al. (2023), these methods perform the steps in one sequence generation process, differing from our definition of compositional reasoning (see Section 2.1).
@@ A.2 Automated Machine learning
Table 4 compares our method with existing methods that employ LLMs for autoML. A detailed analysis of recent advancements in this domain can be found in Tornede et al. (2023). To the best of our knowledge, Text-to-ML is the first method that automatically generates the optimized program of the entire ML program across task modalities.
While recent autoML techniques also include predictor-based methods (Dudziak et al., 2020) and weight-sharing methods (Pham et al., 2018), these techniques are less suitable for the desiderata in our study: applications in low resource constraints scenarios and search space with diverse pretrained deep learning models.
Table 4: Comparison with existing methods in terms of qualitative features. (1) Code generation: whether the method automatically generates programs (2) Deep learning: whether the method automatically generates (part of) programs for deep learning tasks. (3) Data preparation: whether the method automatically chooses the strategy for data preparation (4) Modeling: whether the method automatically chooses the strategy for modeling (5) Optimization: whether the method optimizes the programs (components). * Although the method can automatically choose the strategy for modeling, the method does not automatically generate the program for modeling. ** The method uses the LLM as a pseudo-optimization model instead of a dedicated optimization algorithm (e.g., evolutionary computation and reinforcement learning.)
@@ Methods
GPT-NAS (Yu et al., 2023) $\checkmark$* $\checkmark$
MLCopilot (Zhang et al., 2023a) $\checkmark$*
GENIUS (Zheng et al., 2023) $\checkmark$* $\checkmark$**
Evoprompting (Chen et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$**
CAAFE (Hollmann et al., 2023) $\checkmark$ $\checkmark$
SapientML (Saha et al., 2022) $\checkmark$ $\checkmark$ $\checkmark$
AutoML-GPT (Zhang et al., 2023b) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ *
HuggingGPT (Shen et al., 2023) $\checkmark$ $\checkmark$ $\checkmark$
@@ Appendix B Proofs of Theorems
$=P(Y\in\mathbf{\Omega})$ (1)
Theorem. (Theorem 3.5) For any sequence generation process, if there exists a step in the process that generates the entire output sequence and Assumption 3.4 holds for the step, then the best case complexity of $\mathbb{E}_{x\sim{P}}[G(\Gamma)]$ of the process is $\mathcal{O}(exp(\Gamma))$.
@@
In the best case, $\mathbf{p}_{X}^{\Omega,\lambda}(i)=\epsilon_{max}\;(1\leq i\leq I)$. Then by equation (1), $P(Y\in\mathbb{\Omega})=\epsilon_{max}^{I}$ and $\mathbf{E}(G)=\epsilon_{max}^{-I}$. Therefore, the best case complexity of $\mathbb{E}_{x\sim{P}}[G(\Gamma)]$ is $\mathcal{O}(exp(\Gamma))$.
@@
Then, in the best case, $P(Y\in\Omega)=\epsilon^{\xi_{X}^{\Omega,\lambda}(\epsilon)}$ and $\mathbb{E}_{x\sim{P}}[G(\Gamma)]=\epsilon^{-\xi_{X}^{\Omega,\lambda}(\epsilon)}$.
Theorem. (Theorem 3.8) If Assumption 3.7 holds, then for Contextual Modular Generation as defined in Algorithm 1, the best case complexity of $\mathbb{E}_{x\sim{P}}[G(\Gamma)]$ is $\mathcal{O}(\Gamma)$.
@@
Then by equation (1), in the best case, for each $Y^{j}$, $P(Y^{j}\in\mathbf{\Omega}^{j})=\epsilon_{min}^{Z^{j}}\geq\epsilon_{min}^{M}\;(1\leq j\leq J)$. In Contextual Modular Generation, each module is iteratively generated until being evaluated as valid. Therefore, we have
$\mathbb{E}_{x\sim{P}}[G(\Gamma)]\leq\sum_{j=1}^{\lceil\frac{I}{M}\rceil}\mathbb{E}_{x\sim{P}}[G^{j}{(\Gamma)}]={\lceil\frac{I}{M}\rceil\epsilon_{min}^{-M}}$
Thus, the best case complexity of $\mathbb{E}_{x\sim{P}}[G(\Gamma)]$ is $\mathcal{O}(\Gamma)$.
@@ C.1 Unit Testing Details
1:Input: a progenitor testing protocol $\mathfrak{P}$, an autoregressive LLM $f_{\lambda}$, the textual description of an ML task $\mathbf{T}$, a specification of the targeted module $j$, and the number of versions of synthetic data $N$
@@
4:     Given $\mathbf{T}$ and $\mathfrak{P}$, $f_{\lambda}$ devises a plan $\mathbf{P}$ that details the specific features and data flow of the modules tailored for the task.
5:     Verify $\mathbf{P}$ with $\mathfrak{P}$
6:until $\mathbf{P}$ is allowed by $\mathfrak{P}$
7:With $\mathbf{P}$ and $\mathfrak{P}$, construct the specific unit tests for the ML task $\{\mathbf{E}^{i}\}$
@@
10:         Given $\mathbf{P}$, $f_{\lambda}$ generates program $\mathbf{D}_{n}$, which produces synthetic input and output data for the ML task
11:         Verify $\mathbf{D}_{n}$ with the test for synthetic data in $\{\mathbf{E}^{i}\}$
12:     until $\mathbf{D}_{n}$ passes the test in $\{\mathbf{E}^{i}\}$
@@
14:Given $\mathbf{T}$ and $j$, $f_{\lambda}$ generates the program $\mathbf{C}_{\mathbb{M}}$ for $j$
@@
16:     Verify $\mathbf{C}_{j}$ with $\mathbf{D}_{n}$ and the test for the module in $\{\mathbf{E}^{i}\}$
17:     if $\mathbf{C}_{j}$ passes the test then
@@
Regardless of the specific task and modules, the generated modules should be compatible with a pre-written set of modules for evaluation, training, and testing. Progenitor testing protocols first ensure the external connectivity by testing the form of the input and output data in the ML task, since all modules, including data preparation, modeling, post processing, optimizer, and loss functions, center around the input and output data in the ML task.
@@ C.3 Search Strategy
In our study, we employ two search strategies: random search and BOHB (Falkner et al., 2018). Random search is a powerful baseline in hyperparameter optimization due to its wide applicability and search efficiency for tasks with low effective dimensionality (Bergstra & Bengio, 2012). BOHB combines Bayesian Optimization (BO) and Hyperband (HB) (Li et al., 2018) to optimize hyperparameters efficiently. BOHB inherits HB’s robustness and flexibility, and BO’s ability to ensure strong performance. BOHB starts by rapidly identifying promising configurations like HB and refines them using BO to achieve optimal solutions faster. It’s designed to be simple, computationally efficient, and adaptable to various tasks. For BOHB, we use the default hyperparameters in Falkner et al. (2018) and set 30 epochs as the maximum budget for deep learning tasks.
@@ C.5 ZC proxies details
ZC proxies are an autoML technique that allows for the prediction of a fully trained neural network’s performance with small computational expenses. In this study, we utilize four ZC proxies: flops, params, naswot, and synflow. The term flops stands for Floating Point Operations Per Second, indicating the computational power of the network. Params denotes the total number of parameters within the network, reflecting its complexity. Although flops and params are fundamental concepts within deep learning, they have been validated as powerful baseline ZC proxies (Krishnakumar et al., 2022).
@@ C.7 Hyperparameters for finetuning
Our research focuses on an unusual search space with pretrained models at the architectural level. Öztürk et al. (2022) also explored a similar search space. Consequently, we utilize a comparable set of hyperparameters for fine-tuning the models (Table 5). To tailor this approach to our specific requirements, we have reduced the minimum batch size from 16 to 2 to accommodate the varying GPU memory demands across different tasks.
@@ D.2 Kaggle datasets
Age The ICR - Identifying Age-Related Conditions competition (Kaggle & InVitro Cell Research, 2023) involves the development of machine learning models capable of detecting various age-related conditions from medical imaging data. Participants are provided with a dataset containing images labeled with different age-related conditions, and the task is to create models that can accurately classify these conditions. The competition aims to advance research in medical imaging by leveraging AI to improve the diagnosis and understanding of age-related diseases.
@@
Default The American Express - Default Prediction competition (Kaggle & Express, 2023) challenges participants to predict the likelihood of credit default by American Express cardholders. The dataset provided includes a wide range of anonymized features spanning transactional data, account information, and user behavior over time, designed to simulate real-world conditions. The goal is to enhance credit risk models and improve financial inclusion by accurately identifying at-risk customers.
@@
Whale The Happywhale - Whale and Dolphin Identification competition (Kaggle & Happywhale, 2023) focuses on the identification of individual whales and dolphins from images. Participants are tasked with developing algorithms that can recognize and match cetaceans based on unique physical features captured in photographs, contributing to marine biology research and conservation efforts. The dataset comprises images labeled with species and individual IDs, making it a challenge in both classification and individual identification.
@@
Strip The Mayo Clinic - STRIP AI competition (Kaggle & Clinic, 2023) is aimed at predicting stroke and other thrombotic events by analyzing electroencephalogram (EEG) data. This competition provides a platform for the development of AI models that can assist clinicians in early detection and intervention for stroke patients, leveraging a dataset of EEG recordings with associated clinical outcomes. This challenge stands at the intersection of AI and healthcare, with the potential to significantly impact patient care and outcomes in neurology.
@@
Exam The Kaggle - LLM Science Exam competition (Kaggle, 2023) challenges participants to create language models that can solve science exam questions. The focus is on evaluating the ability of these models to understand and apply scientific knowledge accurately. This competition is part of Kaggle’s efforts to benchmark the performance of language models in educational contexts, providing a unique opportunity to advance AI in the field of automated question answering.
@@
Learning The Feedback Prize - English Language Learning competition (Kaggle & Lab, 2023) aims to develop AI models that can provide feedback on written English tasks by students learning English as a foreign language. The challenge involves analyzing student-written essays to identify areas for improvement, thus aiding in language learning and teaching. This competition seeks to leverage AI to enhance English language education by providing scalable, personalized feedback.
@@ Appendix E Prompts
Figure 15: Prompt for CIFAR-10. The loading code is copied from the official description of the datasets (Krizhevsky et al., 2023)
Figure 16: Prompt for CIFAR-100. The loading code is copied from the official description of the datasets (Krizhevsky et al., 2023)